40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 19 01:49:52 2020, Koji, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 11:27:04 2020, rana, Update, General, M4.5 EQ in LA Sep18-EQ.pdf
       Reply  Sat Sep 19 18:08:34 2020, rana, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 19:14:59 2020, Koji, Update, General, ITMX released / ETMY UR magnet knocked off IMG_6218.jpeg
       Reply  Sun Sep 20 11:41:54 2020, gautam, Update, General, IMC re-locked 
       Reply  Mon Sep 21 12:40:58 2020, gautam, Update, General, Workable IFO recovery IFOrecovery.pngIMG_5345.JPG
Message ID: 15581     Entry time: Sat Sep 19 11:27:04 2020     In reply to: 15580     Reply to this: 15583
Author: rana 
Type: Update 
Category: General 
Subject: M4.5 EQ in LA 

the seismometers obviously saturated during the EQ, but the accelerometers captured some of it. It looks like there's different saturation levels on different sensors.

Also, it seems the mounting of the MC2 accelerometers is not so good. There's some ~10-20 Hz resonance its mount that's showing up. Either its the MC2 chamber legs or the accelerometers are clamped poorly to the MC2 baseplate.

Sun Sep 20 00:02:36 2020 edit: fixed indexing error in plots

* also assuming that the sensors are correctly calibrated in the front end to 1 count = 1 um/s^2 (this is what's used in the summ pages)

Attachment 1: Sep18-EQ.pdf  269 kB  Uploaded Sun Sep 20 01:03:01 2020  | Hide | Hide all
Sep18-EQ.pdf
ELOG V3.1.3-